# Quantum-Based Software Engineering

### v1

@@ Abstract.
Quantum computing has demonstrated potential for solving computationally intensive problems more efficiently than classical methods. Many software engineering tasks, such as test case selection, static analysis, code clone detection, and defect prediction, involve complex optimization, search, or classification, making them candidates for quantum enhancement. In this paper, we propose Quantum-Based Software Engineering (QBSE), a potential research direction for applying quantum computing to classical software engineering problems. We outline its scope, clarify its distinction from quantum software engineering (QSE), and identify key problem types that may benefit from quantum optimization, search, and learning techniques. We also summarize existing research efforts that remain fragmented. Finally, we sketch a preliminary research agenda that may help guide the future development of QBSE as a structured and meaningful direction within software engineering.
@@ 1. Introduction
In this paper, we propose Quantum-Based Software Engineering (QBSE) as a potential research direction for exploring how quantum computing can support classical software engineering tasks. Unlike QSE, which targets quantum software development, QBSE focuses on classical software engineering challenges and investigates whether quantum techniques can improve efficiency or scalability. We suggest that QBSE offers a timely and promising perspective that complements ongoing efforts in quantum software and classical engineering automation. In the following sections, we define the scope of QBSE, outline emerging applications, and outline preliminary ideas for a research agenda.
@@ 2. A Brief Background on Quantum Computing
Quantum computing offers a fundamentally different model of computation, built on the principles of quantum mechanics (nielsen2002quantum, ). This section introduces basic concepts that will help the readers understand the quantum techniques discussed later in this paper.
  • Qubits. Unlike classical bits that are either 0 or 1, a quantum bit (or qubit) can exist in a superposition of both. A qubit’s state, written as $|\varphi\rangle$, can be expressed as a combination of two basis states: $\alpha|0\rangle+\beta|1\rangle$, where $\alpha$ and $\beta$ are complex numbers such that $|\alpha|^{2}+|\beta|^{2}=1$. The values $|\alpha|^{2}$ and $|\beta|^{2}$ represent the probability amplitudes, indicating the likelihood of observing the qubit in state $|0\rangle$ or $|1\rangle$ upon measurement.
  • Quantum Gates and Circuits. Quantum gates manipulate qubits through unitary operations. Common gates include the Hadamard gate (which creates superposition), the Pauli-X gate (analogous to a classical NOT) and the controlled-NOT (CNOT) gate, which is used to create entanglement. A quantum circuit consists of a sequence of such gates applied to one or more qubits.
@@
With these quantum computing foundations in place, we now define what we mean by QBSE and how it differs from related fields.
@@ 3. What is Quantum-Based Software Engineering?
QBSE is conceptually distinct QSE [^1] Alternatively, these two directions (QBSE and QSE) may also be referred to as QC4SE (Quantum Computing for Software Engineering) and SE4QC (Software Engineering for Quantum Computing), respectively. We use the terms QBSE and QSE in this paper for clarity and consistency.. While QSE focuses on the engineering of quantum software systems, including design, programming languages, compilation, testing, and verification (zhao2020quantum, ), QBSE explores how quantum computing can be used to enhance classical software engineering tasks.
@@ 4. Why We Need a New Research Direction
In summary, formally establishing QBSE as a research direction helps consolidate previous efforts, promote collaboration, and accelerate progress in applying quantum computing to classical software engineering.
@@ 5. Quantum Techniques
This section introduces quantum computing techniques that may be useful in solving classical software engineering problems. These techniques are categorized into quantum search, optimization, machine learning, and annealing-based methods that have been explored or proposed in recent work. Table 1 summarizes the main classes of quantum techniques, along with representative algorithms and their hardware requirements. We then briefly review each category in the remainder of this section. While this section focuses on the technical principles of quantum methods, we also briefly mention key software engineering applications for each technique. Detailed task-specific use cases are discussed in Section 6.
@@ 5.2. Quantum Optimization Algorithms
5.3. Annealing-Based Optimization
A typical formulation involves minimizing an Ising Hamiltonian:
$H=\sum_{i}h_{i}\sigma_{i}^{z}+\sum_{i<j}J_{ij}\sigma_{i}^{z}\sigma_{j}^{z},$
@@ 6. Emerging Applications in Quantum-Based Software Engineering
In this section, we present classical software engineering tasks that may benefit from quantum computing techniques. These tasks often involve large search spaces, combinatorial complexity, or high-dimensional data representations, making them suitable candidates for quantum optimization, search, and learning methods. We focus on organizing emerging studies that concretely demonstrate how quantum computing can be applied to classical software engineering problems. We categorize the literature by task, including testing, defect prediction, code analysis, and specification checking, and highlight the quantum techniques explored in each case. While these efforts remain exploratory, they form an initial base for QBSE.
6.1. Test Case Selection and Prioritization
@@ 6.4. Vulnerability Detection and Software Security
Quantum-enhanced models have been investigated for software vulnerability detection, especially in source code embeddings and token classification tasks. QNNs, QSVMs, QNLPs, and quantum embedding models (beer2020training, ; schuld2019quantum, ; zeng2023qnlp, ) have been used to classify vulnerable patterns. For example, Zhou et al. (zhou2022new, ) trained a QNN on tokenized code, and Song et al. (song2024recurrent, ) proposed a recurrent quantum embedded neural network (RQENN) to model context-sensitive patterns.
@@ 6.5. Specification Checking and Component Synthesis
Formal verification tasks, such as property checking and component retrieval, can benefit from a quantum search. Grover-based algorithms have been applied to reachability checking in finite-state machines (hall2009ses, ), including variations that handle unknown target counts. Quantum-enhanced search has also been proposed to synthesize components from libraries based on input-output constraints (hall2009ses, ), although these approaches remain in the proof-of-concept stage and have not yet been validated in realistic settings.
@@ 6.6. Positioning and Summary of Emerging QBSE Applications
In addition to these task-specific studies, several broader efforts have attempted to map quantum algorithm classes to software engineering challenges. Miranskyy et al. (miranskyy2022quantum, ) outlined how eight classes of quantum algorithms, including Grover’s search, quantum SAT solvers, quantum linear systems, and quantum walks, could be applied to various phases of software development. Their work offers a conceptual roadmap, but does not examine specific implementations. Mandal et al. (mandal2025quantum, ) reviewed existing research applying quantum computing to software engineering and categorized it into areas such as optimization, search, and machine learning. Their paper emphasized potential integration workflows, but did not systematically organize empirical studies by task type. Wang et al. (wang2025quantum, ) focused on the role of Quantum Artificial Intelligence (QAI), particularly quantum optimization and quantum machine learning, in software engineering. Their paper outlined challenges such as algorithm design, noise handling, and problem representation, and suggested potential research directions. While this work offers valuable insights into the role of quantum AI within software engineering, its focus represents only one aspect of the broader QBSE landscape. In contrast, our paper provides a more comprehensive view by organizing and analyzing task-specific applications of quantum computing across a wide range of classical software engineering activities.
Compared with these works, our paper aims to establish QBSE as a distinct research direction by systematically organizing and interpreting early applications of quantum computing to classical software engineering tasks. By highlighting the tasks, techniques, and emerging evidence in a structured manner, we seek to define the scope and significance of QBSE within the broader software engineering landscape.
@@ 7. A Preliminary Research Agenda
  • Design of quantum-assisted methods. Once suitable formulations are in place, the next step is to develop quantum-assisted methods that integrate quantum algorithms into classical software engineering workflows. These may include quantum routines for tasks such as test case selection, defect prediction, or code clone detection, potentially embedded within larger software engineering pipelines, though such integrations remain at an early stage.
@@ 8. Conclusion
In this paper, we proposed Quantum-Based Software Engineering (QBSE) as a potential research direction. QBSE focuses on applying quantum computing techniques to solve classical software engineering problems. It differs from QSE, which focuses on developing software for quantum systems.

### v2

@@ Abstract.
Quantum computing has demonstrated the potential to solve computationally intensive problems more efficiently than classical methods. Many software engineering tasks, such as test case selection, static analysis, code clone detection, and defect prediction, involve complex optimization, search, or classification, making them candidates for quantum enhancement. In this paper, we introduce Quantum-Based Software Engineering (QBSE) as a new research direction for applying quantum computing to classical software engineering problems. We outline its scope, clarify its distinction from quantum software engineering (QSE), and identify key problem types that may benefit from quantum optimization, search, and learning techniques. We also summarize existing research efforts that remain fragmented. Finally, we outline a preliminary research agenda that may help guide the future development of QBSE, providing a structured and meaningful direction within software engineering.
@@ 1. Introduction
In this paper, we introduce Quantum-Based Software Engineering (QBSE) as a new research direction to explore how quantum computing can support classical software engineering tasks. Unlike QSE, which targets quantum software development, QBSE focuses on classical software engineering challenges and investigates whether quantum techniques can improve efficiency or scalability. We suggest that QBSE offers a timely and promising perspective that complements ongoing efforts in quantum software and classical engineering automation. In the following sections, we define the scope of QBSE, outline emerging applications, and present preliminary ideas for a research agenda.
@@ 2. A Brief Background on Quantum Computing
Quantum computing offers a fundamentally different model of computation, built on the principles of quantum mechanics (nielsen2002quantum, ). This section introduces basic concepts that will help the reader understand the quantum techniques discussed later in this paper.
  • Qubits. Unlike classical bits that are either 0 or 1, a quantum bit (or qubit) can exist in a superposition of both. The state of a qubit, written as $|\varphi\rangle$, can be expressed as a combination of two basis states: $\alpha|0\rangle+\beta|1\rangle$, where $\alpha$ and $\beta$ are complex numbers such that $|\alpha|^{2}+|\beta|^{2}=1$. The values $|\alpha|^{2}$ and $|\beta|^{2}$ represent the probability amplitudes, indicating the likelihood of observing the qubit in state $|0\rangle$ or $|1\rangle$ upon measurement.
  • Quantum Gates and Circuits. Quantum gates manipulate qubits through unitary operations. Common gates include the Hadamard gate (which creates superposition), the Pauli-X gate (analogous to a classical NOT), and the controlled-NOT (CNOT) gate, which is used to create entanglement. A quantum circuit consists of a sequence of such gates applied to one or more qubits.
@@ 3. What is Quantum-Based Software Engineering?
QBSE is conceptually distinct QSE [^1] In this paper, we use the terms QBSE and QSE for clarity and consistency. For convenience, these two directions can also be informally described as ”quantum computing for software engineering” (QC4SE) and ”software engineering for quantum computing” (SE4QC), respectively, following the conceptual distinction discussed in this paper. While QSE focuses on the engineering of quantum software systems, including design, programming languages, compilation, testing, and verification (zhao2020quantum, ), QBSE explores how quantum computing can be used to enhance classical software engineering tasks.
@@ 5. Quantum Techniques
This section introduces quantum computing techniques that may be useful in solving classical software engineering problems. These techniques are categorized into methods based on quantum search, optimization, machine learning, and annealing that have been explored or proposed in recent work. Table 1 summarizes the main classes of quantum techniques, along with representative algorithms and their hardware requirements. We then briefly review each category in the remainder of this section. While this section focuses on the technical principles of quantum methods, we also briefly mention key software engineering applications for each technique. Detailed task-specific use cases are discussed in Section 6.
@@ 5.3. Quantum Machine Learning
5.4. Annealing-Based Optimization
A typical formulation involves minimizing an Ising Hamiltonian:
$H=\sum_{i}h_{i}\sigma_{i}^{z}+\sum_{i<j}J_{ij}\sigma_{i}^{z}\sigma_{j}^{z},$
@@ 6. Emerging Applications in Quantum-Based Software Engineering
In this section, we present classical software engineering tasks that may benefit from quantum computing techniques. These tasks often involve large search spaces, combinatorial complexity, or high-dimensional data representations, making them suitable candidates for quantum optimization, search, and learning methods. We focus on organizing emerging studies that concretely demonstrate how quantum computing can be applied to classical software engineering problems. We categorize the literature by task, including testing, defect prediction, code analysis, vulnerability detection, and specification checking, and highlight the quantum techniques explored in each case. While these efforts remain exploratory, they form an initial base for QBSE.
6.1. Test Case Optimization and Minimization
@@ 6.4. Vulnerability Detection and Software Security
Quantum-enhanced models have been investigated for software vulnerability detection, especially in source code embeddings and token classification tasks. QNNs, QSVMs, and quantum embedding models (beer2020training, ; schuld2019quantum, ) have been used to classify vulnerable patterns. For example, Zhou et al. (zhou2022new, ) trained a QNN on tokenized code, and Song et al. (song2024recurrent, ) proposed a recurrent quantum embedded neural network (RQENN) to model context-sensitive patterns.
@@ 6.5. Specification Checking and Component Synthesis
Formal verification tasks, such as property checking, can benefit from a quantum search. Grover-based algorithms have been applied to reachability checking in finite-state machines (hall2009ses, ), including variations that handle unknown target counts. Quantum-enhanced search has also been proposed to synthesize components from libraries based on input-output constraints (hall2009ses, ), although these approaches remain in the proof-of-concept stage and have not yet been validated in realistic settings.
@@ 6.6. Positioning and Summary of Emerging QBSE Applications
In addition to these task-specific studies, several broader efforts have attempted to map quantum algorithm classes to software engineering challenges. Miranskyy et al. (miranskyy2022quantum, ) outlined how eight classes of quantum algorithms, including Grover’s search, quantum SAT solvers, quantum linear systems, and quantum walks, could be applied to various phases of software development. Their work offers a conceptual roadmap, but does not examine specific implementations. Mandal et al. (mandal2025quantum, ) reviewed existing research applying quantum computing to software engineering and categorized it into areas such as optimization, search, and machine learning. Their paper emphasized potential integration workflows, but did not systematically organize empirical studies by task type. Wang et al. (wang2025quantum, ) focused on the role of Quantum Artificial Intelligence (QAI), particularly quantum optimization and quantum machine learning, in software engineering. Their paper outlined challenges such as algorithm design, noise handling, and problem representation, and suggested potential research directions. While this work offers valuable insights into the role of quantum AI within software engineering, its focus represents only one aspect of the broader QBSE landscape.
Compared with these works, our paper aims to help establish QBSE as a distinct research direction by systematically organizing and interpreting early applications of quantum computing to classical software engineering tasks. By highlighting the tasks, techniques, and emerging evidence in a structured manner, we seek to define the scope and significance of QBSE within the broader software engineering landscape.
@@ 7. A Preliminary Research Agenda
  • Design of quantum-assisted methods. Once suitable formulations are in place, the next step is to develop quantum-assisted methods that integrate quantum algorithms into classical software engineering workflows. These may include quantum routines for tasks such as test case selection, defect prediction, or code clone detection, potentially embedded within larger software engineering pipelines, although such integrations remain at an early stage.
@@ 8. Conclusion
In this paper, we introduced Quantum-Based Software Engineering (QBSE) as a new research direction. QBSE focuses on applying quantum computing techniques to solve classical software engineering problems. It differs from QSE, which focuses on developing software for quantum systems.